OpenStudy (anonymous):

convert 15 7/8 to a decimal

OpenStudy (anonymous):

15+7/8

OpenStudy (anonymous):

7/8=.875 15+.875=15.875

OpenStudy (anonymous):

7 divided by 8 = 0.875 0.875+15.00= 15.875

OpenStudy (anonymous):

@OrangeMaster is right! :)

OpenStudy (anonymous):

15 + 875= 15.875